# TRELLISWorld: Training-Free World Generation from Object Generators

### v1

@@ 4.3 Quantitative Comparison
Table 1: Perceptual scores with standard deviation and confidence intervals across different methods. Our proposed method achieves the highest CLIP (Radford et al., 2021) and the second-lowest standard deviation, indicating better alignment with users’ prompts.
Method CLIP Mean $\uparrow$ CLIP STD $\downarrow$ 95% Confidence
SynCity 0.260202 0.023223 [0.257264, 0.263140]
inpaint baseline 0.264191 0.025335 [0.260986, 0.267397]
w/o blending 0.262029 0.027042 [0.258608, 0.265450]
TRELLISWorld 0.265201 0.024958 [0.262043, 0.268358]
To compare perceptual alignments with SynCity using CLIP score based on clip-vit-base-patch32 model (Radford et al., 2021), we uniformly rendered $18$ views at close distance from $15$ generation results each of size 4x3x1 across diverse prompts. We adopt the same set of prompts (e.g., ”city”, ”medieval”, ”desert”, …) for both methods following the procedure detailed in subsection A.1. The result in Table 1 only shows marginal improvements as CLIP score does not directly measure blending quality across tiles.
@@
Table 2: Computational resources required to generate a scene. All experiments are conducted on the same $4\times 3$ tile layout, and all $S$ and SLAT samples are generated using $25$ diffusion steps.
Method Total Time Time per Tile VRAM Consumption
SynCity 76 min 46 sec 384 sec $\sim$48 GB
TRELLISWorld 14 min 24 sec 72 sec <<16 GB
All experiments are run on a single NVIDIA GeForce RTX 4090 GPU (48GB). Without any optimization, Table 2 shows that our method achieves a $5\times$ speedup and significantly less memory compared to SynCity (Engstler et al., 2025) for generating tiles of the same size and resolution. Moreover, because our method does not rely on autoregressive inference, larger scenes can potentially be parallelized across multiple GPUs for further speedup.
@@ 5 Applications
Figure 8: City scene expansion results (not cherry-picked) using TRELLISWorld. Given the leftmost 1x1x1 tile as input, the model generates a 3x3x1 extended scene. Three diverse outputs are shown to the right, demonstrating variations.
@@ 6 Limitations
As a training-free approach, our method is inherently constrained by the capabilities of the underlying base models. In particular, the performance of TRELLIS directly limits both the visual fidelity and efficiency of our scene generation pipeline.
@@ A.1 Text Prompt
Our method gives the user explicit control of the 3D prompts used in generation. Below is an example of 3D prompt we use to generate our $4\times 3\times 1$ city theme:
@@
For other themes (“city”, “medieval”, “desert”, “cyberpunk”, “ancient Rome”, “minecraft”, “forest”, “ocean”, “winter”, “lego”, “park”, “amusement park”, “airport”, “college”, “room”), we ask LLMs to generate similar prompts using in-context learning from the city prompt above.
@@ A.2 Inpaint
Our method can fill missing regions or extend a user-provided tile. We use RePaint (Lugmayr et al., 2022) with a Gaussian-blurred mask to partially preserve the edges of the input tile. This encourages smoother transitions between the generated and existing content.
A.3 Additional Qualitative Results
@@
Figure 11: Additional (not cherry-picked) 4x3x1 Tiled Gaussian Splatting Generation Results: city, desert, medieval.
Figure 12: Additional (not cherry-picked) 4x3x1 Tiled Gaussian Splatting Generation Results: ocean, room, winter.
A.4 Disclosure

### v2

@@ 4.3 Quantitative Comparison
Table 1: Perceptual scores and CannyAvg error with standard deviation and confidence intervals across different methods. Our proposed method achieves the highest CLIP (clip) score and lowest CannyAvg error, indicating better alignment with users’ prompts as well as minimal seams between tiles.
Method CLIP Mean $\uparrow$ CLIP STD CLIP 95%
SynCity 0.26020 0.02322 [0.25726, 0.26314]
inpaint baseline 0.26419 0.02534 [0.26099, 0.26740]
avg. blending 0.26203 0.02704 [0.25861, 0.26545]
TRELLISWorld 0.26520 0.02496 [0.26204, 0.26836]
CannyAvg Mean $\uparrow$ CannyAvg STD CannyAvg 95%
SynCity 7.81725 2.31850 [6.98759, 8.64691]
inpaint baseline 7.71797 5.63712 [5.70075, 9.73519]
avg. blending 6.55861 6.32146 [4.29650, 8.82072]
TRELLISWorld 5.61331 5.49558 [3.64674, 7.57987]
To compare perceptual alignments with SynCity using the CLIP score based on clip-vit-base-patch32 model (clip), we uniformly rendered $18$ views at close distance from $15$ generation results, each of size 4x3x1 across diverse prompts, visualized in Figure 14. We adopt the same set of prompts (e.g., ”city”, ”medieval”, ”desert”, …) for both methods following the procedure detailed in subsection A.1. The result in Table 1 only shows marginal improvements, as the CLIP score does not directly measure blending quality across tiles.
Seam Visibility
The main contribution of our method lies in reducing seams between incoherent object generations. To measure the visibility of the seams, we first rendered the scene using orthographic projection and then calculated a binary map using the Canny edge detection operator. We use the average intensity of this binary map to define the seam visibility metric CannyAvg. You can find a detailed explanation of CannyAvg in subsubsection A.3.2.
551010151522334455Stride $s$ ($S=16$)CannyAvg(a) CannyAvg by Different Strides
05,0005{,}00010,00010{,}00022334455Avg. Sec./Scene(b) CannyAvg by Computation
Figure 8: CannyAvg measuring seams in a scene using variable stride $s$. All experiments are conducted on a 4x3x1 chunk layout with a fixed prompt. Decreasing stride increases scene quality, with $s=8$ being an optimal trade-off between computational time and scene quality.
@@
05510101515202025250200200400400600600Number of Chunks in the SceneAvg. Sec./ChunkRuntime by Different Number of ChunksSynCityTRELLISWorld
Figure 9: Computational cost required to generate a scene of variable scene sizes. All experiments are conducted on square chunk layouts from 1x1x1 to 5x5x1 with a fixed prompt. TRELLISWorld takes less time to generate a scene for all scene sizes.
551010151505,0005{,}00010,00010{,} [^0] Stride $s$ ($S=16$)Avg. Sec./Scene(a) Runtime by Different Strides
020020040040005,0005{,}00010,00010{,} [^0] Number of Tiles(b) Runtime by Different Tiles
Figure 10: Computational cost required to generate a scene for variable stride $s$. All experiments are conducted on the 4x3x1 chunk layouts with a fixed prompt. Decreasing $s$ reciprocally increases the number of tiles, but the time to generate a tile remains constant ($27.21$ seconds per tile on average)
Without any optimization, TRELLISWorld takes $77.96$ seconds on average to generate a chunk, which is a $5.80\times$ speedup over SynCity (syncity) ($452.04$ seconds). TRELLISWorld also significantly reduces memory cost. During comparison to SynCity, we found that SynCity is unable to run on a single NVIDIA GeForce RTX 4080 (16GB). Therefore, we resort all SynCity experiments to a single NVIDIA GeForce RTX 4090 (48GB), a better GPU on all metrics, giving SynCity an unfair advantage. Figure 9 and Figure 10 show a linear increase in runtime as the number of chunks or tiles increases. Moreover, because our method does not rely on autoregressive inference, larger scenes can potentially be parallelized across multiple GPUs for further speedup.
@@ 5 Applications
Figure 11: City scene expansion results (not cherry-picked) using TRELLISWorld. Given the leftmost 1x1x1 chunk as input, the model generates a 3x3x1 extended scene. Three diverse outputs are shown to the right, demonstrating variations.
@@ 6 Limitations
As a training-free approach, our method is inherently constrained by the capabilities of the underlying base models. In particular, the performance of TRELLIS directly limits both the visual fidelity and the efficiency of our scene generation pipeline.
@@ 7 Conclusion
Acknowledgement
We sincerely thank Zekai Gu for sharing resources, and anonymous reviewers for their valuable feedback.
@@ A.1 Text Prompt
Our method gives the user explicit control of the 3D prompts used in generation. Below is an example of a 3D prompt we use to generate our 4x3x1 city theme:
@@
For other themes (“city”, “medieval”, “desert”, “cyberpunk”, “ancient Rome”, “minecraft”, “forest”, “ocean”, “winter”, “lego”, “park”, “amusement park”, “airport”, “college”, “room”), we ask LLMs to generate similar prompts using in-context learning from the city prompt above. For example, we generated our medieval prompt by asking ”[IN_CONTEXT_PROMPT] Above is a prompt for creating a large city block with parks, please give a prompt for ’a medieval tile with farmland fields and cottages’ following the same format.”
As shown above, the input prompt is formatted as a 3D tensor. Therefore, prompts in spatial proximity will remain close in the generated scene. We do not force the user to provide a tensor that has the exact shape of our tiles, as this may limit flexibility and ease of use. Instead, we treat the input prompt tensor as a tensor spanning the entire scene. Specifically, given a tile location, we sample the nearest prompt in the prompt tensor. This way, users can provide a coarser prompt tensor.
@@ A.2 Inpaint
Our method can fill missing regions or extend a user-provided chunks. We use RePaint (inpaint) with a Gaussian-blurred mask to partially preserve the edges of the input chunks. This encourages smoother transitions between the generated and existing contents.
A.3 Metrics
A.3.1 CLIP
We measured CLIP similarity using $18$ views per generated scene. Figure 14 shows $6$ different views from SynCity and TRELLISWorld respectively. We used the mega-prompt (e.g., ’a medieval tile with farmland fields and cottages’) instead of the object-level prompt for CLIP text input, appending ”, brid eye view” and removing all auxiliary modifications such as ”, on top of a base” and ”, square tile” used by SynCity and TRELLISWorld to better reflect the actual content in the images.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
SynCity
TRELLISWorld
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 14: Sampled views for CLIP similarity calculation.
A.3.2 CannyAvg
We used CannyAvg to evaluate the amount of seams in our generated results. The CannyAvg value for each scene is calculated as the average pixel intensity of an orthographic top-down rendering after Canny edge detection with a $[200,400]$ threshold. For each 4x3x1 generated result, we rendered an image at 1536x2048 resolution. To show the effectiveness of CannyAvg, we have included Canny edge detection results in Figure 15, along with other possible edge detection methods that are less effective at highlighting seams.
s=2
Refer to caption Refer to caption Refer to caption Refer to caption
s=16
Refer to caption Refer to caption Refer to caption Refer to caption
RGB Canny Laplacian Sobel
Figure 15: Visualization of different edge detection methods. Canny best aligns with human perception of seams, while other methods highlight internal content, making them unsuitable for quantitative seam measurement.
A.4 Extension to Other Base Generator
To demonstrate the extensiveness of our method, we have tested our method using two other base generators. The first is an unconditional diffusion transformer trained on Minecraft terrains and the second is an image-conditioned triplane diffusion transformer trained on crops of Minecraft structures. While both architectures only support generation of 16x16x16 blocks layout, we adapt them using our method to generate structures with sizes beyond what they were trained on. The qualitative results in Figure 16 shows that our method keeps the generation cohesive without any visible seams.
(a) Terrain Generator
(b) Structure Generator
Figure 16: Generation results of tiled diffusion applied to two different Minecraft object generators. (a) The Minecraft terrain generator is trained on 16x16x16 blocks but sampled at 32x32x16 blocks. (b) The Minecraft structure generator is trained on 16x16x16 blocks but sampled at 22x21x14 blocks.
A.5 Extension to Image-Conditioned Generator
We argue that our method is best applied to a text-conditioned object generator. Applying our method to image-conditioned object generators often result in unaligned floor level, which is an indicator of unsuccessful blending. We show generation results using the image-conditioned TRELLIS model in Figure 17.
We suspect that this discrepancy between text-conditioned and image-conditioned model lies in the difference in their underlying learned conditional distributions of the score models $\theta_{img},\theta_{text}$. When the model $\theta_{img},\theta_{text}$ is trained on the same set of 3D object but with different conditional label $c_{img},c_{text}$. They have the same marginalized distribution $q_{\theta_{img}}(x_{0})=\int q_{\theta_{img}}(x_{0}\mid c_{img})p(c_{img})dc_{img}=\int q_{\theta_{text}}(x_{0}\mid c_{text})p(c_{text})dc_{text}=q_{\theta_{text}}(x_{0})$. However, their conditional distribution is different: $q_{\theta_{text}}(x_{0}\mid c_{text}=\text{specific-text})\neq q_{\theta_{img}}(x_{0}\mid c_{img}=\text{specific-img})$ even when the given ”specific-text” and ”specific-img” are associated with the same 3D object. Specifically, $q_{\theta_{text}}(x_{0}\mid c_{text})$ might be more diffuse compared to $q_{\theta_{img}}(x_{0}\mid c_{img})$, which makes the text-conditioned model more robust to out-of-data-distribution queries of the score function.
City
Refer to caption Refer to caption Refer to caption
Minecraft
Refer to caption Refer to caption Refer to caption
Village
Refer to caption Refer to caption Refer to caption
Image Left Image Right Generated Result
Figure 17: Visualization of applying TRELLISWorld to image-conditioned generator. We generated 2x1x1 chunk which was split to 3 tiles (left, right, and middle tile) from 2 conditional images. The middle image was chosen arbitrarily from left or right images. In all generations give inconsistent floor levels. The ”Village” example shows significant artifact.
A.6 Additional Qualitative Results
@@
Figure 18: Additional (not cherry-picked) 4x3x1 chunk Tiled Gaussian Splatting Generation Results: city, desert, medieval.
Figure 19: Additional (not cherry-picked) 4x3x1 chunk Tiled Gaussian Splatting Generation Results: ocean, room, winter.
A.7 Disclosure